If the rms speed of the molecules of a quantity of helium at temperature 142 K is the same as the rms speed of neon atoms at a second temperature, what is the temperature of the neon gas? Answer in Kelvins. The answer is NOT 710, 713.55 or 284.
Answer #1

The rms speed of a gas is given by

v_{rms} = \sqrt{\frac{3 k_{B}T}{m}} = \sqrt{\frac{3 RT}{M}}

where kB is Boltzmann's Constant, T is temprature in Kelvin R is gas constant = 8.31 JK-1mol-1 and M is molecular mass,

Now For Helium M = 4

v_{rms}(He) = \sqrt{\frac{3 \times 8.31\times 142}{4}}

Now For Neon M = 20

v_{rms}(Ne) = \sqrt{\frac{3 \times 8.31\times T}{20}}

From the data given

v_{rms}(Ne) = v_{rms}(He)

\sqrt{\frac{3 \times 8.31\times T}{20.1797}} = \sqrt{\frac{3 \times 8.31\times 142}{4.00}}

\frac{3 \times 8.31\times T}{20.1797} = \frac{3 \times 8.31\times 142}{4}

T = \frac{20.1797 \times 142}{ 4} = 716.379 K \approx 716.4 K